5 years ago, Barton Industries issued 25-year noncallable, semiannual bonds with a $1,000 face value and a 6% coupon, semiannual payment ($30 payment every 6 months). The bonds currently sell for $894.87. If the firm's marginal tax rate is 25%, what is the firm's after-tax cost of debt? Do not round intermediate calculations. Round your answer to two decimal places. %
Answer #1

The after tax cost of debt is calculated by first computing the before cost of debt.

The yield to maturity is calculated to compute the before cost of debt.

Information provided:

Face value= future value= $1,000

Market price= present value= $894.87

Time= 25 years - 5 years= 20 years *2= 40 semi-annual periods

Coupon rate= 6%/2= 3%

Coupon payment= 0.03*1,000= $30

The yield to maturity is calculated by entering the below in a financial calculator:

FV= 1,000

PV= -894.87

N= 40

PMT= 30

Press the CPT key and I/Y to compute the yield to maturity.

The value obtained is 3.4917

Therefore, the yield to maturity is 3.4917%*2= 6.9833% 6.98%.

After tax cost of debt= before cost of debt*(1 - tax rate)

= 6.9833%*(1 - 0.25)

= 5.2375% 5.24%.

In case of any query, kindly comment on the solution.
Answer #2

SOLUTION :


Using online calculator for YTM ,  ref : https://dqydj.com/bond-yield-to-maturity-calculator/ :


Key -in the following :


Current market price ($) = 894.87


Bond Face Value ($) = 1000


Years to maturity (25 - 5) = 20 


Annual coupon rate (%) = 6.0


Frequency of coupon = 2 (being semi-annual) 



And see the result : YTM = Pre tax cost of debt = 6.983 %  


So, 


After tax cost of debt = 6.983( 1 - 0.25) = 5.23725 = 5.24 % (ANSWER)

answered by: Tulsiram Garg